OpenStudy (anonymous):

can someone please simplify sqrt3(-80)^4?

OpenStudy (anonymous):

Is it a cube root or a square root?

OpenStudy (anonymous):

70944801.078

OpenStudy (mathmale):

I think you mean "the third root of (-80)^4" The third root is not a sqrt. You could present math problems more clearly using either Equation Editor or Draw (below).\[\sqrt[3]{(-80)^4}\]

OpenStudy (mathmale):

Please clarify your meaning before others get involved in helping you.

OpenStudy (mathmale):

\[\sqrt[3]{(-80)^4}\] can be re-written (for easier evaluation) as:\[\sqrt[3]{(-80)^4}=\sqrt[3]{(-80)^3}\sqrt[3]{(-80)}\]

OpenStudy (mathmale):

and \[\sqrt[3]{(-80)}\]can be further separated as follows:\[\sqrt[3]{-8}*\sqrt[3]{10}\]

OpenStudy (anonymous):

That is the third root of negative eight raised to the fourth power

OpenStudy (mathmale):

But what about the 10, as in -80 = -8 * 10?

OpenStudy (anonymous):

I have choices. I am to simplify it and most of them have to do with an imaginary number

OpenStudy (mathmale):

Do you mean (third root of -8)^4 or do you mean third root of (-8)^4? I know this situation is probably new for you, but we have spent a lot of time on this problem without being really clear about what is meant. Again, please consider using Equation Editor or Draw. Have you used Draw before?

OpenStudy (mathmale):